OpenStudy (anonymous):

Help with (1/4x)-(3/4)=7 Find X.

OpenStudy (anonymous):

Adding both sides by 3/4 would be a good start.

OpenStudy (anonymous):

(1-4x)-(3/4)=7 Add both sides by 3/4. (1-4x) = 7+(3/4) What's next?

OpenStudy (anonymous):

It's a rational equation. I have do it like this (4/4)*(1/4x)-(4x/4x)*(3/4)=7

OpenStudy (anonymous):

Wait, so you actually mean (1/4x)-(3/4)=7, not (1-4x)-(3/4)=7?

OpenStudy (anonymous):

Yes I messed up my question :/

OpenStudy (anonymous):

Well, this doesn't change the first step. Just add both sides by 3/4 to make it easier to find x.

OpenStudy (anonymous):

You can do it, right?

OpenStudy (anonymous):

But it's not the correct way to do it. I have to get the lcd

OpenStudy (anonymous):

Are you sure the question is to find the value of x in (1/4x)-(3/4)=7?

OpenStudy (agent0smith):

\[\frac{ 1 }{ 4x }-\frac { 3} {4}=7\] Multiply both sides by 4x like so: \[4x \left( \frac{ 1 }{ 4x }-\frac { 3} {4}\right)= 4x \times 7\]distribute on the left \[\frac{ 1 \times 4x }{ 4x }- \frac { 3 \times 4x } {4}= 7 \times 4x \]

OpenStudy (anonymous):

Yeah. I don't see how this has to do with LCD or LCM...

OpenStudy (agent0smith):

I guess if you wanted to find the LCD you could multiply the (3/4) by x/x... \[ \frac{ 1 }{ 4x }-\frac { 3} {4} \times \frac{ x }{ x }= 7\]to get\[\frac{ 1 }{ 4x }-\frac { 3x} {4x}= 7\]then you can add the fractions with common denominator 4x\[\frac{ 1 -3x}{ 4x }= 7\]

OpenStudy (anonymous):

Thanks that's what I need!

OpenStudy (agent0smith):

You're welcome :)